OpenStudy (erinweeks):

Find the vertical asymptotes, if any, of the graph of the rational function.

OpenStudy (marissalovescats):

Vertical Asymptotes are what make the denominator equal zero.

OpenStudy (erinweeks):

\[f(x)=\frac{ x-4 }{ x(x-4) }\]

OpenStudy (erinweeks):

Can you help me with this.. show me how to work it out

OpenStudy (anonymous):

which values of x makes this function undefined? [in this case it's when the denominator = 0]

OpenStudy (erinweeks):

0 and 4??

OpenStudy (marissalovescats):

I'm pretty sure that's what it is :)

OpenStudy (erinweeks):

these are my possibilites A. x = 4 and x = 4 B. x = 4 C. x = 0 and x = 4 D. no vertical asymptote

OpenStudy (anonymous):

you are right. it's C

OpenStudy (erinweeks):

thank you

OpenStudy (marissalovescats):

Good job

OpenStudy (anonymous):

actually

OpenStudy (marissalovescats):

@jim_thompson5910 Says hello Vertical Asymptotes.

OpenStudy (erinweeks):

am i wrong?

OpenStudy (anonymous):

this answer is only 0 i think @ErinWeeks @marissalovescats since this function = 1/x

OpenStudy (erinweeks):

there is no only zero thats why i chose C. because i think 4 would work also

OpenStudy (anonymous):

unless the numerator is written wrong. 4 does not work

OpenStudy (anonymous):

4 also works

OpenStudy (anonymous):

0/0=undefined doesn't it?

OpenStudy (anonymous):

if you graph it, x=4 says error and x=0 also says error, so there are vertical asymptotes at these to x values

OpenStudy (erinweeks):

so i was correct?

OpenStudy (anonymous):

its obviously an error, but I don't believe it's a physical vertical asymptote on a graph

OpenStudy (anonymous):

techincally you are i guess; especially since it's the only option. it mathematically an asymptote but I don't think it is graphically

OpenStudy (anonymous):

true